A ladder is resting against a wall. The top of the ladder touches the wall at a height of 18 feet. Find the length of the ladder if the length is 6 feet more than its distance from the wall
Could someone help me with this?
Thank you

Answer by richwmiller(17219)

The wall and the floor form a right angle.
The hypotenuse is the ladder
a^2+b^2=c^2
a=18 height
b=c-6 distance from wall
If the length means the length of the ladder
"if the length is 6 feet more than its distance from the wall"
18^2+(c-6)^2=c^2
18^2+c^2-12c+36=c^2
18^2-12c+36=0
18*18+18*2-12c=0
18(18+2)-12c=0
18*20=12c
3*20=2c
30=c
The ladder is 30 ft
The foot of the ladder is 24 feet from the wall
check
18^2+24^2=30^2
ok
Just curious, what did you think would happen if you didn't ask "Could someone help me with this?"?